# The ML Supply Chain in the Era of Software 2.0: Lessons Learned from Hugging Face

### v1

@@ Abstract.
The last decade has seen widespread adoption of Machine Learning (ML) components in software systems. This has occurred in nearly every domain, from natural language processing to computer vision. These ML components range from relatively simple neural networks to complex and resource-intensive large language models. However, despite this widespread adoption, little is known about the supply chain relationships that produce these models, which can have implications for compliance and security. In this work, we conduct an extensive analysis of 760,460 models and 175,000 datasets mined from the popular model-sharing site Hugging Face. First, we evaluate the current state of documentation in the Hugging Face supply chain, report real-world examples of shortcomings, and offer actionable suggestions for improvement. Next, we analyze the underlying structure of the extant supply chain. Finally, we explore the current licensing landscape against what was reported in prior work and discuss the unique challenges posed in this domain. Our results motivate multiple research avenues, including the need for better license management for ML models/datasets, better support for model documentation, and automated inconsistency checking and validation. We make our research infrastructure and dataset available to facilitate future research.
@@ 1. Introduction
The use of machine learning (ML) models in software applications has increased dramatically over the last decade, including in recommendation systems, computer vision, chatbots, image generation, automated software engineering, and more. As creating and training new models has become increasingly more expensive (Patterson et al., 2021), developers have turned to fine-tuning pre-existing models. This approach may save time and effort, but it also introduces the complexity of a new ML supply chain, complete with many novel challenges.
@@
While supply chains for conventional software typically consist of software components, libraries, configuration files, and processes (Tan et al., 2022; Gao et al., 2024), for ML-intensive systems, the supply chain is more complex, comprising all elements that make an ML algorithm work. Creating a modern ML system involves the integration of ML models and traditional software components, as well as different types of reuse or integration, taking into account dependencies between models themselves or between models and their training datasets. In addition, the training of ML models can rely on multiple datasets or dataset aggregates, and datasets, in turn, each have their own supply chain (Lee et al., 2023).
@@
Adaptation or reuse of existing models can also introduce degrees of complexity. Models can be fine-tuned (Church et al., 2021) or quantized from existing models (Gholami et al., 2022). (Fine-tuning allows a developer to use an existing model to create a model that specializes in certain tasks or to change the model’s behavior, such as removing safety or ethical protections. Quantization compresses an existing model such that it can be run on smaller hardware with minimal performance loss.) Different ML models can be combined to form a single architecture that can, in turn, be reused or adapted (Ardabili et al., 2019). Finally, the outputs from one model can be used to train another model, such as through synthetic data (Marwala et al., 2023) or the use of a teacher-student approach (Hinton, 2015).
@@
The ML supply chain, particularly those involving generative models (Lee et al., 2023), is not necessarily linear, progressing cleanly from one dependency or stage to the next. Instead, as outlined by Lee et al. (Lee et al., 2023), there can be branches and even cycles in the supply chain. For example, datasets can be used not only in the initial training of models but also for the fine-tuning of pre-existing base models, sometimes in multiple instances and by different developers. Additionally, some stages of the supply chain can backfeed into others, making relationships increasingly complex, if not recursive. For example, the outputs of a generative model can be added to pre-existing datasets and used for training future versions. The sheer volume of content from generative models makes this situation increasingly likely, if not inevitable.
@@
While previous work has considered the documentation (Pepe et al., 2024), evolution (Jiang et al., 2022), and environmental impact (Castaño et al., 2024a) of ML models, the complexity and challenges of the ML supply chain as a whole have not yet been fully explored. The supply chain relationships between models and dependent GitHub repositories have been explored by Pepe et al. (Pepe et al., 2024) and Jiang et al. (Jiang et al., 2024b), but the relationships between ML models remain understudied.
To contribute to bridging this gap, we aim to investigate the emerging ML model supply chain on Hugging Face (Inc., 2024) (hereafter referred to as the ‘ML supply chain’), specifically the documentation practices of model owners, the structure and complexity of the supply chain itself, and the existence and prevalence of potential compliance issues. We note that proper documentation and documentation practices are foundational to mapping, and thus managing the ML supply chain. Launched in 2016, Hugging Face is, as of January 2025, the largest repository of ML models and datasets (Jiang et al., 2023c, 2022). As with GitHub and traditional software, Hugging Face provides a central hub for model developers and data scientists to explore, share, and experiment with ML models. While other model forges—such as TensorFlow Hub (Google Inc., 2024), PyTorch Hub (The Linux Foundation, 2024), and Model Zoo (Koh, 2024)—exist, the Hugging Face platform has the widest reach, hosting more than 750K models and 175K datasets across many different use cases. Hugging Face therefore represents a robust opportunity to understand ML supply chains and analyze how developers interact with them.
We mine models and datasets hosted on Hugging Face, using the available information to construct a supply chain graph. We highlight different types of documentation inconsistencies that complicate the creation of this supply chain graph, discuss the underlying structure and features of the current ML supply chain on Hugging Face, and analyze the present licensing landscape. Ultimately, we demonstrate that there is still much work to be done to facilitate better dependency/supply chain management on Hugging Face and in the ML ecosystem.
@@
  • An analysis of the interactions between models and datasets in the supply chain.
@@
2. Background and Related Work
@@
The ML supply chain has emerged as a complex structure with many steps and components, each of which can affect the final output of a given model. Lee et al. identify eight stages in the generative AI supply chain, but most apply more broadly: 1) the creation of expressive works that will eventually be used to train a model, 2) the conversion of these expressive works into digital data, 3) the compilation of these data points into training datasets, 4) the creation of an ML model by selecting an architecture, training datasets, and a training algorithm, 5) the fine-tuning of existing “base” models, 6) the deployment of the trained/fine-tuned model as a service, 7) the use of the model to generate output, and finally 8) applying additional alignment to further improve the model or meet user needs (Lee et al., 2023). Modern ML models are also becoming increasingly complex with respect to their architectures. For example, Mixture of Expert (MoE) models and ensemble models comprise multiple specialized models trained by dividing the problem space into homogeneous regions (Masoudnia and Ebrahimpour, 2014; team, 2023). Recent work has also proposed using smaller ML models to replace the hidden states within a Recurrent Neural Network (RNN) architecture to facilitate better recall and compression of context (Sun et al., 2024). Both of these examples demonstrate that understanding the relationships among models and among model architectures will become increasingly difficult.
@@
As a result, sharing information about models has become critically important. ML components do not, however, always disclose data sources, making it difficult to comply with or even keep track of licensing obligations associated with the data (Hassan et al., 2024). This creates conditions that could lead to legal disputes over the use of copyrighted material in training ML components (Vincent, 2022). Model cards (Mitchell et al., 2019) have become the standard method for sharing information about ML models hosted on Hugging Face (Face, ndb), including intended uses, limitations, and datasets used in training (Face, nda). Although tools are in development to streamline the process (Face, ndc), model card creation on Hugging Face is primarily a manual process in which a user enters documentation about a model into a predefined markdown template (Face, 2024b). This template is robust, including spaces for model description, uses, bias, limitations, testing, etc., as well as optional fields for citation, technical specifications, and environmental impact. The manual nature of documentation, however, introduces a significant potential for human error, ambiguity, and incompleteness (Jiang et al., 2024b, 2023d). Indeed, the quality and adoption of model cards remains low (Bhat et al., 2023), despite user studies and other efforts by Hugging Face (Face, [n. d.]) to ameliorate the situation, and despite researchers proposing tools such as DocML (Bhat et al., 2023) to create and evolve model cards.
Hugging Face has previously been mined to study the maintenance and evolution of ML models (Castaño et al., 2024b), security risks in the ML supply chain (Jiang et al., 2022), the carbon footprint of ML models (Castaño et al., 2023), and the documentation of ML models in terms of datasets, bias, and licenses (Pepe et al., 2024). Through an analysis of discussion forums and a mining study of Hugging Face, Taraghi et al. (Taraghi et al., 2024) identified challenges and benefits associated with model reuse. Also, previous work identified challenges and deficiencies in the organization of the ML supply chain, including poorly chosen or erroneous model names (Jiang et al., 2023a) and frequently missing or incorrect licensing information (Longpre et al., 2023) on Hugging Face and other platforms.
@@
While prior work has previously noted the incompleteness of model/dataset documentation on Hugging Face (Jiang et al., 2023c; Pepe et al., 2024; Yang et al., 2024), our work goes beyond this observation and provides a detailed analysis of specific examples in which documentation is incomplete and erroneous. These real-world observations can help model owners avoid common documentation pitfalls, empower model hubs to implement automated checks for common errors, and make researchers aware of potential noise in data mined from Hugging Face.
Oreamuno et al. (Oreamuno et al., 2024) also observed that the majority of models and datasets on Hugging Face lacked documentation. When present, this documentation was in the form of model/data cards, was often incomplete, and followed standards inconsistently. Whereas the work of Oreamuno et al. focused on exploring the presence and completeness of documentation, we report various inconsistencies and anomalies that are introduced by Hugging Face users during the process of documenting their models and datasets. The anomalies we encountered can exist even in cases where the model/dataset documentation is complete.
Yang et al. (Yang et al., 2024) conducted an in-depth analysis of the state of dataset documentation on Hugging Face, focusing primarily on the content included in the documentation (e.g., description, structure, creation). Our study expands on this work by also considering the documentation of ML models in addition to datasets and by noting errors or inconsistencies introduced by model/dataset owners.
Jiang et al. (Jiang et al., 2024a) studied naming practices and naming anomalies for models on Hugging Face. They found that model names carry more semantic information than do the names of traditional software packages, so anomalies, such as misrepresenting the model architecture, can lead to greater confusion in the ML context. Our work does not consider the accuracy of pre-trained model (PTM) names when compared against the metadata they provide; rather we note difficulties in constructing the ML supply chain given naming inconsistencies and ambiguities in metadata references.
Pepe et al. (Pepe et al., 2024) conducted the study most closely related to ours. However, instead of focusing on the structure of the ML supply chain, they explore the extent to which models document datasets, potential biases, and licenses. Both Pepe et al. (Pepe et al., 2024) and Jiang et al. (Jiang et al., 2024b) analyze a supply chain, but they focus on the relationship between models and downstream GitHub repositories that make use of those models. We instead focus on the relationships among PTMs, dependent models, and datasets.
@@
Both Pepe et al. (Pepe et al., 2024) and Jiang et al. (Jiang et al., 2024b) also explore model licensing on Hugging Face. In the case of Jiang et al. (Jiang et al., 2024b), this is done as a short demonstration illustrating the utility of the PEATMOSS dataset. We extend these prior works by collecting and analyzing more recent data, which reflects the current state of model/dataset licensing. Both works also consider the interplay of model licenses with the licenses of dependent GitHub repositories. Neither work, however, evaluates license incompatibilities between models or between models and datasets. We further expand on prior work by considering various license classes, multi-licensing, and the failure to comply with license-imposed naming requirements. Where previous work examines the relationships between projects and models, we examine the relationships between multiple models. Thus, our work is useful not only to software developers, as is the case for prior work, but also to data scientists, those who want to reuse models that have complex dependencies, and those who want to develop new models. By providing an ecosystem-level examination of the ML supply chain, we provide these groups with the information necessary to understand how models relate to each other and evolve, licensing concerns regarding ML components on Hugging Face, and potential challenges that might be encountered regarding model documentation.

### v2

# An Empirical Analysis of Machine Learning Model and Dataset Documentation, Supply Chain, and Licensing Challenges on Hugging Face

@@ Abstract.
The last decade has seen widespread adoption of Machine Learning (ML) components in software systems. This has occurred in nearly every domain, from natural language processing to computer vision. These ML components range from relatively simple neural networks to complex and resource-intensive large language models. However, despite this widespread adoption, little is known about the supply chain relationships that produce these models, which can have implications for compliance and security. In this work, we conducted an extensive analysis of 760,460 models and 175,000 datasets extracted from the popular model-sharing site Hugging Face. First, we evaluate the current state of documentation in the Hugging Face supply chain, report real-world examples of shortcomings, and offer actionable suggestions for improvement. Next, we analyze the underlying structure of the existing supply chain. Finally, we explore the current licensing landscape against what was reported in previous work and discuss the unique challenges posed in this domain. Our results motivate multiple research avenues, including the need for better license management for ML models/datasets, better support for model documentation, and automated inconsistency checking and validation. We make our research infrastructure and dataset available to facilitate future research.
@@ 1. Introduction
The use of machine learning (ML) models in software applications has increased dramatically over the last decade, including in recommendation systems, computer vision, chatbots, image generation, and automated software engineering. As creating and training new models has become increasingly more expensive (Patterson et al., 2021), developers have turned to fine-tuning pre-existing models. This approach may save time and effort, but it also introduces the complexity of a new ML supply chain, complete with many novel challenges.
@@
Supply chains for conventional software typically consist of software components, libraries, configuration files, and processes that contribute to the software (Williams et al., 2025; Okafor et al., 2022; Wang et al., 2025). However, for ML-intensive systems, the typical ML supply chain is more complex, comprising all elements that make an ML algorithm work (Tan et al., 2022; Gao et al., 2024). Creating a modern ML system involves the integration of ML models and traditional software components, taking into account dependencies between models themselves or between models and their training datasets. In addition, the training of ML models can rely on multiple datasets or dataset aggregates, and datasets, in turn, each have their own supply chain. Because proper license compliance requires developers to consider the permissions granted and requirements imposed by any individual component’s license, the complexity of ML supply chains raises new and challenging issues, particularly given that many questions remain regarding the applicability of copyright law to generative AI (Lee et al., 2023a).
@@
Adaptation or reuse of existing models can also introduce degrees of complexity (Davis et al., 2023). Models can be fine-tuned (Church et al., 2021) or quantized from existing models (Gholami et al., 2022). (Fine-tuning allows a developer to use an existing model to create a model that specializes in certain tasks or to change the model’s behavior, such as removing safety or ethical protections. Quantization compresses an existing model such that it can be run on smaller hardware with minimal performance loss.) Different ML models can be combined to form a single architecture that can, in turn, be reused or adapted (Yang et al., 2024b). Finally, the outputs from one model can be used to train another model, such as through synthetic data (Marwala et al., 2023) or the use of distillation (Hinton, 2015). (Distillation is a technique where a smaller, simpler student model is trained to mimic the behavior of a larger, more complex teacher model.)
@@
The ML supply chain, particularly those involving generative models (Lee et al., 2023a), is not necessarily linear, progressing cleanly from one dependency or stage to the next. Instead, as outlined by Lee et al. (Lee et al., 2023a), there can be branches and even cycles in the supply chain. For example, datasets can be used not only in the initial training of models but also for the fine-tuning of pre-existing base models, sometimes in multiple instances and by different developers. Additionally, some stages of the supply chain can backfeed into others, making relationships increasingly complex, if not recursive. For example, the outputs of a generative model can be added to pre-existing datasets and used for training future versions. The sheer volume of content from generative models makes this situation increasingly likely, if not inevitable.
@@
While previous work has considered the documentation (Pepe et al., 2024), evolution (Jiang et al., 2022), and environmental impact (Castaño et al., 2024a) of ML models, the complexity and challenges of the ML supply chain as a whole have not yet been fully explored. The supply chain relationships between models and dependent GitHub repositories have been explored by Pepe et al. (Pepe et al., 2024) and Jiang et al. (Jiang et al., 2024c), but the relationships between ML models remain understudied.
To contribute to bridging this gap, we aim to investigate the emerging ML model supply chain on Hugging Face (Inc., 2024) (hereafter referred to as the “ML supply chain”), specifically the documentation practices of model owners, the structure and complexity of the supply chain itself, and the existence and prevalence of potential license compliance issues. Launched in 2016, Hugging Face is, as of September 2025, the largest repository of ML models and datasets (Jiang et al., 2023c, 2022). As with GitHub and traditional software, Hugging Face provides a central hub for model developers and data scientists to explore, share, and experiment with ML models. While other model forges—such as TensorFlow Hub (Google Inc., 2024), PyTorch Hub (The Linux Foundation, 2024), and Model Zoo (Koh, 2024)—exist, the Hugging Face platform has the widest reach, hosting more than 750K models and 175K datasets across many different use cases. Hugging Face therefore represents a robust opportunity to understand ML supply chains and analyze how developers interact with them.
We mine models and datasets hosted on Hugging Face, using the available information to construct a supply chain graph. We highlight different types of documentation inconsistencies that complicate the creation of this supply chain graph, discuss the underlying structure and features of the current ML supply chain on Hugging Face, and analyze the present licensing landscape. We note that proper documentation and documentation practices are foundational to mapping, and thus managing, the ML supply chain. Ultimately, we demonstrate that there is still much work to be done to facilitate better dependency/supply chain management on Hugging Face and in the ML ecosystem.
@@
  • An analysis of the interactions between models and datasets in the ML supply chain on Hugging Face.
@@
  • An in-depth manual analysis of the top models and datasets on Hugging Face.
@@
2. Background
In this section, we provide the background necessary to contextualize and understand our results. Specifically, we highlight the ML supply chain, software licensing, and documentation practices for ML components on Hugging Face.
2.1. The ML supply chain
The ML supply chain has become a complex structure with many steps and components, each of which can affect the final output of a given model. Lee et al. (Lee et al., 2023a) identify eight events in the generative AI supply chain, but most apply to AI systems more broadly:
  1. (1)
    The creation of expressive works that will eventually be used to train a model;
  2. (2)
    The conversion of these expressive works into digital data;
  3. (3)
    The compilation of these data points into training datasets;
  4. (4)
    The creation of an ML model by selecting an architecture, training datasets, and a training algorithm;
  5. (5)
    The fine-tuning of existing “base” models;
  6. (6)
    The deployment of the trained/fine-tuned model as a service;
  7. (7)
    The use of the model to generate output, and finally
  8. (8)
    Applying additional alignment to further improve the model or meet user needs.
More generally, the ML supply chain comprises various components, including datasets, models, and traditional software. These components relate to each other in several key ways, including data examples being collected to create a dataset, multiple datasets being aggregated to create a composite dataset, datasets being used to train ML models, models being fine-tuned on pre-trained base models, and models being incorporated into software projects. These complex relationships distinguish the ML supply chain from other software supply chains: different steps in the process of creating and deploying ML models can feed back into each other in various, nonlinear ways, as identified by Lee et al. (Lee et al., 2023a), and different types of dependency relationships can relate to or build off of the original works in different ways, as identified by Duan et al. (Duan et al., 2024). Understanding how work relates is important, as reusing any component through a supply chain can raise security risks or licensing implications (Wang et al., 2025).
@@ 2.1. The ML supply chain
Modern ML models are also becoming increasingly complex with respect to their architectures. For example, Mixture of Expert (MoE) models and ensemble models comprise multiple specialized models trained by dividing the problem space into homogeneous regions (Masoudnia and Ebrahimpour, 2014; team, 2023). Recent work has also proposed using smaller ML models to replace the hidden states within a Recurrent Neural Network (RNN) architecture to facilitate better recall and compression of context (Sun et al., 2024). Both examples demonstrate that understanding the relationships among models and model architectures will become increasingly difficult.
Hugging Face, an online platform that hosts ML components, including models and datasets, has become a central hub for those who use and create ML components, showing exponential growth in recent years (Castaño et al., 2024b). Users can download components provided by others and upload their own components. It has become the leading registry for pre-trained models online (Jones et al., 2024), presenting new opportunities for reusing ML components (Davis et al., 2023). By studying the Hugging Face environment, we aim to uncover how multiple ML components can relate to one another and consider the implications of this critical supply chain.
2.2. Software Licensing
In the same way as other creative works, software is protected by copyright law in the United States (cop, 2021).A copyright holder has several rights under U.S. copyright law, including the right to reproduce the work, create derivative works, distribute the work, and authorize others to engage in these activities (USC, 2021). Although the copyright owner can transfer some or all of the rights in a work, an owner who wishes to retain copyright while authorizing use by another party in a way that would otherwise be infringing typically does so by means of a license, which allows the copyright owner to impose conditions on the use (Meeker, 2017). In practice, developers often select pre-existing licenses drafted by other organizations and entities rather than drafting new licenses (Almeida et al., 2019; Wintersgill et al., 2024). This means that a particular license may or may not be a good fit for the subject of the license, the intended uses, or the copyright status of the underlying work.
The licensing landscape for ML models, particularly in Hugging Face, is very diverse, including open source licenses (such as GPL and MIT) and Creative Commons licenses, as well as new ML-specific licenses such as Open-RAIL and Llama. These licenses broadly fall into two categories: permissive and restrictive (sometimes called copyleft). Permissive licenses (such as MIT (mit, [n. d.]), Apache-2.0 (apa, [n. d.]), and CC-BY-4.0 (cc-by-4.0, [n. d.])) typically impose few restrictions, often only requiring the user to provide notice files containing attribution or license information. On the other hand, copyleft licenses (such as GPL (gpl, [n. d.]) and CC-BY-SA-4.0 (cc-by-sa-4.0, [n. d.])) typically require that derivative works be made available under the same or compatible licensing (leading some to characterize such licenses as “viral”).
ML-specific licenses follow a similar pattern, but often also impose additional restrictions on model usage, typically involving ethical issues, limiting commercial use beyond a certain number of active users, and prohibitions on using model output to train competing models (Moran, 2021; Aguilera, 2025). These additional restrictions have led some to consider many ML-specific licenses as falling outside the open-source definition (Wiggers, 2025). In response, more new licenses—such as OpenMDW (Foundation, 2025; Gatto, 2025)—have emerged, implementing the Model Openness Framework (White et al., 2024), which aims to align model development with open-source principles.
License compliance is a task with which developers often struggle (Vendome et al., 2018). These challenges are likely heightened in the ML supply chain, given the uncertainties and ambiguities involved in interpreting OSS and CC licenses in the new ML context (Duan et al., 2024; Hardy, [n. d.]). Developers frequently treat license compliance as a lower priority, partly due to limited enforcement (Wintersgill et al., 2024). However, non-compliance can lead to financial (Vance, 2010; Gross, 2007), legal (Claburn, 2022; Brittain, 2023; Andersen v. Stability AI Ltd., [n. d.]; Authors Guild v. OpenAI Inc., 2023), and reputational (Claburn, 2023; Gunningham et al., 2004) consequences for developers and organizations, particularly organizations of sufficient size and prominence to attract attention for non-compliance. More specifically, to the extent that a licensed work is protected by copyright law, use beyond the scope of the license may constitute copyright infringement, which can result in litigation by the licensor, demands for payment or other remedies, and/or undesirable publicity that affects the licensee’s standing in the community.
2.3. ML component documentation
Comprehensive component documentation is essential to understanding the ML supply chain and maintaining license compliance. This need has become increasingly critical as the complexity of the components and the processes involved in their creation and deployment continue to grow. However, ML components do not always disclose data sources, making it difficult to comply with or keep track of all licensing obligations associated with each component (Hassan et al., 2024). This can create conditions that lead to legal disputes over the use of copyrighted material in training ML models (Vincent, 2022), including not only the use of copyrighted works for training without permission but also the use of the generations of one model to train another model, an activity that may not have been permissible under that model’s licensing terms (Meta, 2024).
@@ 2.3. ML component documentation
Model cards (Mitchell et al., 2019) have become the standard method for sharing information about ML models hosted on Hugging Face (Face, ndb), including intended uses, limitations, and datasets used in training (Face, nda). Although tools are in development to streamline the process (Face, ndc), model card creation on Hugging Face is primarily a manual process in which a user enters documentation about a model into a predefined markdown template (Face, 2024b). This template is robust, including fields for model description, uses, bias, limitations, testing, and optional fields for citation, technical specifications, and environmental impact. However, the manual nature of documentation introduces a significant potential for human error, ambiguity, and incompleteness (Jiang et al., 2024c, 2023d). In fact, the quality and adoption of model cards remain low (Bhat et al., 2023), despite user studies and other efforts by Hugging Face (Face, [n. d.]) to improve the situation, and despite researchers proposing tools such as DocML (Bhat et al., 2023) to create and evolve model cards.
3. Related work
Hugging Face has been studied in recent years to investigate the maintenance and evolution of ML models (Castaño et al., 2024b), security risks in the ML supply chain (Jiang et al., 2022), the carbon footprint of ML models (Castaño et al., 2023), and the documentation of ML datasets and models (Pepe et al., 2024; Oreamuno et al., 2024; Longpre et al., 2023).
3.1. Model and dataset documentation
Previous research has identified several documentation shortcomings on Hugging Face. Works by Oreamuno et al. and Jiang et al. highlight that the majority of models and datasets hosted on Hugging Face lack documentation and that less-popular components are more likely to have incomplete documentation (Oreamuno et al., 2024; Jiang et al., 2023c). Additional studies have explored the fields most likely to be missing from the dataset documentation (Yang et al., 2024a), the presence and quality of the license information (Longpre et al., 2023; Pepe et al., 2024), and the documentation of known model biases (Pepe et al., 2024). In addition, some studies have considered other deficiencies, such as poorly chosen or erroneous model names (e.g., nonadherence to conventional naming schemes that lead to confusion) (Jiang et al., 2023a, 2024b).
@@ 3.1. Model and dataset documentation
We build on these previous analyses by identifying specific and previously unreported documentation failures, such as component reference ambiguities, encountered in our analysis of the ML supply chain on Hugging Face (Section 5). Our detailed analysis and enumeration of specific, concrete examples of incomplete and erroneous documentation go beyond previous work. These real-world observations can help model owners avoid common documentation pitfalls, empower model hubs to implement automated checks for common errors, and make researchers aware of potential noise in data mined from Hugging Face.
3.2. The ML supply chain
By analyzing discussion forums and a mining study of Hugging Face, Taraghi et al. (Taraghi et al., 2024) identified various challenges and benefits associated with model reuse. However, they did not explicitly examine the dependency relationships that result from reuse. Other studies, such as those by Pepe et al. (Pepe et al., 2024) and Jiang et al. (Jiang et al., 2024c), have primarily focused on the relationship between models and downstream GitHub repositories that utilize those models. In contrast, our work instead focuses on the relationships among pre-trained/base models,dependent models, and datasets. While Jones et al. do consider model-to-model relationships, they stop short of constructing a comprehensive supply chain graph (Jones et al., 2024). We extend this line of research by constructing such a graph and conducting a robust statistical analysis of its structure. We also conduct a detailed manual review of the most popular models (Section 8) and datasets hosted on Hugging Face (Section 9).
3.3. Software Licensing
@@
Much previous work has investigated software licensing for traditional software, including software licensing issues (Vendome et al., 2018; Kapitsaki et al., 2020; Wu et al., 2017) and practitioners’ understanding of software licensing (Almeida et al., 2019; Wintersgill et al., 2024). More recently, some work has investigated the licensing issues that arise when ML components are used by open-source projects. Works by Pepe et al. (Pepe et al., 2024) and Jiang et al. (Jiang et al., 2024c) explore model licensing on Hugging Face, primarily reporting the presence of licensing information as well as the usage frequencies for those licenses. In the case of Jiang et al. (Jiang et al., 2024c), this was done as a short demonstration illustrating the utility of their PEATMOSS dataset. Both works also consider the interplay of model licenses with the licenses of dependent GitHub repositories. Neither work, however, evaluates license incompatibilities between models or between models and datasets, as we do in this work. We extend these prior works by collecting and analyzing more recent data, which reflects the current state of model/dataset licensing.
The work of Duan et al. (Duan et al., 2024) proposes a tool (ModelGo) and framework for identifying potential license compliance issues involving ML components but provides only artificial examples and offers no empirical analysis of the prevalence of the potential problems in practice. Instead, we consider real-world examples of potential license conflicts and offer a rich, empirical analysis, including a report of license usage statistics, an examination of the frequencies of license differences between parent and children models, and thorough case-studies on a sample of the top models and datasets from Hugging Face. We further expand on prior work by considering various license classes, multi-licensing, and failing to comply with license-imposed naming requirements.
3.4. Summary